Paint costing $18 per gallon is to be mixed with $2 per gallon thinner to make 48 gallons of a paint that can be sold for $14 per gallon. How much paint thinner should be used?
Answer by Nate(3500)

(18(x) + 2(48 - x))/(48) = 14
18x + 2(48 - x) = 672
18x + 96 - 2x = 672
16x = 576
x = 36
48 - 36 = 12 gallons of thinner